RHIC Running: Good running over the weekend. Peak luminosities of 25 x 10^28 cm^-1 s^-1. have been achieved.
There was a kink in the beam lifetimes which was due to the stochastic cooling tripping off. It was fixed early Saturday, improving the beam lifetime. Additional troubles with the stochastic cooling kicker drive. RF experts had to be called over the weekend because the kicker for the Yellow longitudinal stochastic cooling suffered a hardware failure. The same failure occurred this morning between fills.
There was also a fix to the program used to compute the live orbit corrections. This fix went in midday on Sunday.
There was a discussion comparing the amount of Au beam that makes it through transition from the 2007 and 2008 runs. We are operating with ~10% fewer bunches, and ~10% less intensity per bunch. Attempts to put more beam in the machine lead to instabilities.
Sparking in LINAC Tank 3 has caused the BLIP intensity to be reduced by half. They get turned off tonight at midnight to allow work on the Tank. Hopefully it will be recovered during this week's Maintenance Day.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |